OpenStudy (rational):

find all \(n\) such that \(17 | (8n^5-10)\)

OpenStudy (rational):

@mathmath333 lets use below table and solve this http://gyazo.com/444240ce9febfabbb76292d99d1a499a

OpenStudy (rational):

yes that table is using the primitive root 3 it has all indices for all the integers less than 13

OpenStudy (mathmath333):

this primitive roots stuff is above my level

OpenStudy (rational):

thats okay, you're doing very good for an 1 hour of knowing them :)

OpenStudy (mathmath333):

lol

OpenStudy (rational):

i feel you can work this problem with whatever you know so far

OpenStudy (rational):

\[8n^5\equiv 10 \pmod{17}\]

OpenStudy (rational):

take log both sides (ind)

OpenStudy (rational):

\[\text{ind}~ (8n^5) \equiv \text{ind} ~10 \pmod{17}\]

OpenStudy (rational):

use the properties and expand the keeping in mind that your goal is to solve "n"

OpenStudy (rational):

\[\text{ind}~ (8n^5) \equiv \text{ind} ~10 \pmod{17}\] use product property : \[\text{ind}~ 8 + \text{ind}~n^5 \equiv \text{ind} ~10 \pmod{17}\]

OpenStudy (rational):

next use exponent property \[\text{ind}~ 8 + 5*\text{ind}~n \equiv \text{ind} ~10 \pmod{17}\]

OpenStudy (rational):

from the table find out ind 8 and ind 10 plug them in above

OpenStudy (mathmath333):

u can write \(ind ~8=3ind ~2\) ?

OpenStudy (rational):

ofcourse yes, but since we have the table with index values for all numbers, we can simply use them... from the table ind 8 = 10, right ?

OpenStudy (mathmath333):

if u dont have a table then how can one solve them

OpenStudy (rational):

we can create that table or we can solve it using the trick which you have found just now

OpenStudy (rational):

ind 8 = 3*ind 2 = 3*14 = 42 = 10 (mod 16)

OpenStudy (rational):

as you know creating that table is not that hard, for now we may simply use that table..

OpenStudy (mathmath333):

it is coming \(\text{ind} ~n\equiv 2~mod~17\) ?

OpenStudy (rational):

lets see

OpenStudy (rational):

there is a typo earlier, let me fix it quick : \[\text{ind}~ 8 + 5*\text{ind}~n \equiv \text{ind} ~10 \pmod{\color{Red}{16}}\]

OpenStudy (rational):

now work it

OpenStudy (mathmath333):

how comes \( \color{Red}{16}\)

OpenStudy (rational):

the indices properties work in mod\(p-1\) 16 = 17-1

OpenStudy (mathmath333):

ok i see

OpenStudy (rational):

\[\text{ind}~ 8 + 5*\text{ind}~n \equiv \text{ind} ~10 \pmod{\color{Red}{16}}\] from the table ind 8 = 10 ind 10 = 3 plugging them in we have : \[10+ 5*\text{ind}~n \equiv 3 \pmod{\color{Red}{16}}\]

OpenStudy (rational):

subtract 10 both sides and see if you can solve "ind n"

OpenStudy (rational):

if it helps replace "ind n" by \(x\) for time being : \[10+ 5*x \equiv 3 \pmod{\color{Red}{16}}\] solve \(x\)

OpenStudy (mathmath333):

\( 5*x \equiv 9 \pmod{\color{Red}{16}}\)

OpenStudy (rational):

Yes! solve \(x\)

OpenStudy (mathmath333):

m stuck

OpenStudy (rational):

find a value of \(x\) such that \(5x-9\) is divisible by \(16\)

OpenStudy (mathmath333):

wait \(x=21\)

OpenStudy (rational):

we're in mod 16, so..

OpenStudy (rational):

21 is same as 5 in mod 16 so x = 5 is the solution in mod 16

OpenStudy (rational):

\[5*x \equiv 9 \pmod{\color{Red}{16}}\] solving you get \[x \equiv 5 \pmod{\color{Red}{16}}\] but we let ind n = x earlier : \[\text{ind} ~n = 5\]

OpenStudy (rational):

look at the table and see which number has index as 5

OpenStudy (mathmath333):

5

OpenStudy (rational):

Yes so the final solution is \[n \equiv 5\pmod{17}\]

OpenStudy (rational):

that means all the integers of form \(5 + 17k\) where \(k\) is any integer, satisfy the given division problem

OpenStudy (rational):

plugin n=5 and see if 8n^5-10 is really divisible by 17

OpenStudy (mathmath333):

ohk ,

OpenStudy (mathmath333):

yes it is

OpenStudy (rational):

once you practice 1-2 problems, you will see why these work..

OpenStudy (rational):

also this is a one line problem once you get familiar with the method

OpenStudy (mathmath333):

i can be familier with that only if i practice it by myself

OpenStudy (mathmath333):

seems like u r a ghost of ganeshie8

OpenStudy (rational):

lol okay here few more problems for you to practice whenever you're free or want to http://gyazo.com/07dd55c512d6a68a2097af4949bb8a8e

OpenStudy (rational):

you can tag me or ganeshie anytime :)

OpenStudy (mathmath333):

ohk thnx :) for problems

OpenStudy (mathmath333):

ohk lol

Nnesha (nnesha):

you are right @mathmate333 ghost i guess something wrong but dk

OpenStudy (anonymous):

:'( cool problem